OpenStudy (meilendurcer):

1.The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

\[d = ut + 0.5at^2\]subtract ut from both sides\[d - ut = 0.5at^2\]divide both sides by 0.5 (this is the same as multiplying by 2)\[2(d - ut) = at^2\]divide both sides by t^2\[\frac{ 2d - 2ut }{ t^2 } = a\]done ^_^

OpenStudy (meilendurcer):

wow thanks!

OpenStudy (meilendurcer):

@Euler271 this is kind of a 5 part thing, can you help me?

OpenStudy (anonymous):

most likely. what else is there?

OpenStudy (meilendurcer):

2. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

\[d = 0.5g t^2\] divide both sides by 0.5t^2\[\frac{ 2d }{ t^2 } = g\] they want the value of g if t = 10 and d = 490 \[g = \frac{ 2(490) }{ 10^2 } = 9.8\] the units is feet/s^2 . this is not on earth lol ^_^

OpenStudy (meilendurcer):

haha definitely not. 3 more parts if you can help.

OpenStudy (anonymous):

ya. bring it on. please try them first though ^_^

OpenStudy (meilendurcer):

I do but then I realize I could never truly get through it all without help lol. 3.Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

divide both sides by w2 \[d_2 = \frac{ w_1 \times d_1 }{ w_2 }\]

OpenStudy (meilendurcer):

w2 or w^2?

OpenStudy (anonymous):

w_2

OpenStudy (anonymous):

well. w2. lol

OpenStudy (meilendurcer):

oh ok ha it should be acceptable either way

OpenStudy (meilendurcer):

4.Marcie wants to enclose her yard with a fence. Her yard is in the shape of a triangle attached to a rectangle. See the figure below. (I'll insert in a min) The area of this figure can be found by the formula A = (wh) + 0.5(bh). If Marcie wants the total area to be larger than a specified value, she can use the formula A > (wh)+ 0.5(bh). Rewrite this formula to solve for b. Show all steps in your work.

OpenStudy (anonymous):

\[A > wh + 0.5bh\]subtract wh from both sides\[A - wh > 0.5bh\]divide both sides by 0.5h\[\frac{ 2A - 2wh }{ h } > b\]same as\[b < \frac{ 2A - 2wh }{ h }\]

OpenStudy (anonymous):

same as\[b < \frac{ 2A }{ h } - 2w\]

OpenStudy (meilendurcer):

it's not showing up

OpenStudy (meilendurcer):

can you try re-typing that??

OpenStudy (anonymous):

i don't think that would help it show up. try: refreshing page restarting internet browser (copy/paste link at top so you don't have to find it again)

OpenStudy (meilendurcer):

I did all that but thanks! thanks for the help!!

OpenStudy (anonymous):

\[b < \frac{ 2A−2wh }{ h }\]or\[b < \frac{ 2A }{ h } - 2w\]